OpenStudy (anonymous):

1) Describe how to transform (6 sq. rt. x^5)^7 into an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (cwrw238):

the square root of something is same as something to the power (1/2) so sqrt x^5 = (x^5)^1/2 can you continue?

OpenStudy (anonymous):

No, I'm really confused... :/ can you explain what you mean to me?

OpenStudy (cwrw238):

well for example sqrt4 = 4^1/2 = 2 dare you familiar with the laws of indices?

OpenStudy (cwrw238):

* are

OpenStudy (anonymous):

That made a lot more sense! I don't have a clear idea of how to use them, but I know what they are.

OpenStudy (cwrw238):

if we take (x^5) ^ 1/2 it equals x^ (5 * 1/2)

OpenStudy (cwrw238):

= x^(5/2) (6 sq. rt. x^5)^7 = 6^7 * x^(5/2)^7 can you continue?